### v1

@@ Abstract
Recent advancements in generative AI have suggested that by taking visual prompt, GPT-4V can demonstrate significant proficiency in image recognition task. Despite its impressive capabilities, the financial cost associated with GPT-4V’s inference presents a substantial barrier for its wide use. To address this challenge, our work introduces Collage Prompting, a budget-friendly prompting approach that concatenates multiple images into a single visual input. With collage prompt, GPT-4V is able to perform image recognition on several images simultaneously. Based on the observation that the accuracy of GPT-4V’s image recognition varies significantly with the order of images within the collage prompt, our method further learns to optimize the arrangement of images for maximum recognition accuracy. A graph predictor is trained to indicate the accuracy of each collage prompt, then we propose an optimization method to navigate the search space of possible image arrangements. Experiment results across various datasets demonstrate the cost-efficiency score of collage prompt is much larger than standard prompt. Additionally, collage prompt with learned arrangement achieves clearly better accuracy than collage prompt with random arrangement in GPT-4V’s visual recognition.
Keywords:
GPT-4V Large Language Models Efficient Inference
@@ 1 Introduction
With the rapid development of generative AI, various large language models (LLMs) [4, 45, 29] have emerged as generative tools. Beyond text, these models have expanded their capabilities to include text-to-image generation such as Stable Diffusion [25], and text-to-video generation, as seen in Sora [21]. ChatGPT [3], as the most well-known LLMs, has shown it can have natural and coherent conversations, making it a powerful tool in daily life and different industry fields. As the latest version of ChatGPT, GPT-4V is a multi-modal LLM capable of processing both text and images. This capability allows it to be applied to a wider range of applications and tasks. There are many technical reports and user studies [16, 17, 27, 35, 41, 47] about GPT-4V, which conducted thorough evaluations of its capabilities from various aspects.
@@
In [38], the visual capabilities of GPT-4V are investigated within the framework of zero-shod visual recognition tasks, such as image and video recognition. The evaluation of visual capabilities is quite straightforward: images and candidate categories are directly fed into GPT-4V for relevance ranking, yielding Top-1 and Top-5 prediction results. Video and point cloud data are uniformly sampled to generate a set of images, which are then processed by GPT-4V for visual recognition. GPT-4V has achieved remarkable performance across various visual recognition tasks, surpassing previous customized solutions[34, 10]. However, its financial cost associated with its inference can be significant. Specifically, performing image recognition on the ImageNet-1K dataset  [26] requires approximately \$1 for every 20 images, leading to a total evaluation cost of over \$2,500 for the entire dataset. If we consider the rate limites of maximum API requests per minute, the costs will be even higher. Thus, employing GPT-4V for visual recognition tasks is expensive, and it is meaningful to adopt a more budget-friendly way for GPT-4V’s visual recognition.
@@
Figure 1: Visual recognition of GPT-4V with different prompting ways. (a): Standard prompting takes one image as visual prompt for each GPT-4V’s run. (b): Collage prompting concatenates multiple images into one visual prompt and predicts class for all images in each inference. (c): The arrangement of images within collage prompting leads to significantly different results. Green indicates an accurate prediction while red indicates an wrong prediction.
@@
In GPT-4V’s visual recognition, it takes only a single image as the visual prompt. This standard prompting way fails to fully release the potential capacity of GPT-4V, which is able to process multiple inquiries within one prompt simultaneously. Motivated by this idea, we propose Collage Prompting, a new visual prompting way for GPT-4’s visual recognition for budget-friendly inference. In collage prompt, multiple images are collaged into one visual prompt with equal size. GPT-4V is then requested for recognizing the class for all images within this prompt. The overall process is shown in Figure 1b. Since collage prompting way allows for the recognition of multiple images in one run of GPT-4V, it significantly reduces the overall cost. Furthermore, based on the observation in Figure 1c that different arrangements of images within collage prompt could lead to rather large variance of accuracy in GPT-4V’s recognition, we propose to Learn to Collage Prompt (LCP) based on the idea of genetic algorithm. The collage prompt is represented as a graph and a collage predictor is used for estimating the expected accuracy of this collage prompt. LCP is then used for searching for the best arrangement in several iterations.
The contributions of this paper can be summarized as follows,
  • We propose a budget-friendly prompting approach for GPT-4V. By involving multiple images into a single visual prompt, GPT-4V can process multiple images in an inference run, thus reducing the overall expense greatly.
  • We collect the evaluation dataset of collage prompt. The datasets contain various collage prompts from the ImageNet-1K training set and their accuracy in GPT-4V’s image recognition. This dataset is meaningful for studying the effectiveness of collage prompting.
  • We propose a genetic algorithm-based optimization method for collage arrangement. This approach aims to find the optimal arrangement for collage prompts and maximize accuracy in image recognition within GPT-4V.
@@ 2.1 Exploration of GPT-4V
The state-of-the-art large multi-modal model GPT-4V was firstly launched at September 2023 and has demonstrated its strong visual capability in different fields. Early works [39, 41] conducted user study of GPT-4V, where operations were completed by entering prompts on a web interface for GPT-4V. Furthermore, the release of GPT-4V API in November 2023 opened up new opportunities for both academic community and industry to thoroughly evaluate GPT-4V’s performance across various visual benchmarks and provide quantitative data beyond what user studies can offer. GPT-4V’s capacities in multimodal medical diagnosis has been explored in [36, 42, 9], where GPT-4V can process different imaging modalities like CT and MRI in medical scene. GPT-4V can also be utilized to operate robots from providing instructions by taking multimodal input in autonomous driving [8, 12, 35] and task planning [31, 32, 14]. Additionally, GPT-4V has been widely used in advancing video understanding [17], conducting OCR recognition [27], acting as an intelligent web agent [46], and dealing with each observation data [43].  [38] is the first work that considers to adopt extensive quantitative analysis utilizing the established visual benchmarks. However, the evaluation of GPT-4V on visual benchmarks could lead to large expense, since every inference made by GPT-4V will cost money. Hence, it is important to adopt a budget-friendly inference scheme for the evaluation of GPT-4V.
2.2 Prompt Engineering in LLMs
@@
Prompt engineering has emerged as a crucial technique for unlocking the potential of pre-trained large language models (LLMs) and vision-language models (VLMs). The concept of prompt engineering was initially explored and popularized in the LLMs [18, 30, 5] and VLMs [37, 1]. The most common prompting way is zero-shot prompting [24, 6], which offers a paradigm shift in leveraging large LLMs. This method significantly reduces the dependency on vast amounts of training data by employing strategically formulated prompts to steer the model towards executing new tasks. While the primary focus in the field has been on creating prompts that can release the potential capacities of LLMs, we focus on developing a prompting approach that prioritizes cost-efficiency..
3 Methodology
GPT-4V has enabled us to perform comprehensive visual recognition. However, each inference performed by GPT-4V incurs a financial cost, which is determined by the number and type of input and generated tokens [^1] https://openai.com/pricing. Specifically, for image recognition tasks involving images of $512\times 512$ resolution, approximately 5000 tokens are consumed per image. Standard prompting of GPT-4V involves presenting a single image as a visual prompt and processing each image in the dataset individually. With this standard prompting, the expense of evaluating a dataset with 10,000 images could exceed \$500. This prompting way is expensive, and a more budget-friendly way is to process multiple images simultaneously in a single inference run.
Motivated by this idea, we propose Collage Prompting, an efficient alternative to standard visual prompting. Collage Prompting involves concatenating multiple images into a single visual prompt, allowing for simultaneous processing in a single inference run. For example, employing a nine-grid collage prompting approach can decrease expenses to just $1/9$ of what standard individual image prompting incurs. Moreover, given that there are $9!=362,880$ possible configurations for arranging images within a nine-grid collage prompt, identifying the optimal arrangement to maximize GPT-4V’s visual recognition accuracy becomes a crucial problem. And we aim to find the best configuration that ensures high accuracy in GPT-4V’s visual recognition tasks.
3.1 Collage Prompting
GPT-4V’s basic workflow for visual recognition involves taking visual samples as prompts and accompanying them with a textual description that enumerates the categories present in the dataset, as shown in Figure 1a. This combination of visual and text prompts enables GPT-4V to match the visual input with the correct categories. The GPT-4V model analyzes the image, categorizes it based on relevance to the provided text prompt, and predicts the class of the visual input. The effectiveness of GPT-4V in this context is often measured by its Top-1 accuracy in image recognition.
@@ 3.1 Collage Prompting
Figure 2: (a): The workflow of forming the collage prompt from a set of images and related indexes. For a set of images X with two different position indexes I, we can obtain two collage prompts $\textbf{M}_{1}$ and $\textbf{M}_{2}$. Regarding X as the node of a graph, the adjacency matrix of $\textbf{M}_{1}$ and $\textbf{M}_{2}$ can be represented as $\textbf{A}_{1}$ and $\textbf{A}_{2}$. (b): The average accuracy of collage prompts within evaluation datasets using ‘Worst’, ‘Average’ and ‘Best’ arrangement.
When applying GPT-4V to image recognition tasks across datasets with $N$ visual samples using the standard prompting method, it necessitates $N$ separate inferences. Considering the cost associated with each inference, the total expense of evaluating an entire dataset can be substantial. To mitigate these costs, we propose the concept of collage prompting, which allows for the evaluation of multiple visual samples within a single inference run. By assembling $K$ images into one comprehensive visual prompt, we can significantly decrease the inference cost to $1/K$ of the standard method. The collage prompt M is designed to be a $\sqrt{K}\times\sqrt{K}$ grid and each grid contains one image. For example, a collage of four images might be presented in a quadrant grid, while nine images could be arranged in a nine-grid format.
@@
Supposing we have a set of $K$ images $\textbf{X}=[x_{1},x_{2},...,x_{K}]$ and its related position indexes $\textbf{I}=[i_{1},i_{2},...,i_{K}]$, where $i_{j}$ indicates the position number of image $x_{j}$ in the collage prompt, starting from 00 in the top left corner to $K-1$ in the bottom right corner. Hence, the row position $r$ and column position $c$ of image $x_{j}$ can be specified as $r=\lfloor(i_{j}-1)/\sqrt{K}\rfloor$ and $c=(i_{j}-1)\mod\sqrt{K}$ respectively. While collage prompt M has the same size as the standard prompt, GPT-4V can thus take this collage prompt as input and generate the predicted class for all images within the collage prompt.
3.1.1 Collage Prompt as Graph.
Regarding the collage prompt M as graph $\textbf{M}=(\textbf{A},\textbf{F})$ with $K$ nodes, each node $f_{i}\in\mathbb{R}^{l}$ denotes the feature of $x_{i}$, which is extracted from CLIP’s encoder [23]. The adjacency matrix $\textbf{A}\in\mathbb{R}^{K\times K}$ suggests the relative positions of X, where two adjacent images are considered to have an undirected edge. For two images $x_{p}$ and $x_{q}$ that satisfies either $|c_{p}-c_{q}|=1$ when $r_{p}=r_{q}$ or $|r_{p}-r_{q}|=1$ when $c_{p}=c_{q}$, we consider these two images have an edge and set $\textbf{A}[p,q]=\textbf{A}[q,p]=1$. The workflow of representing the collage prompt as a graph is illustrated in Figure 1(a).
3.2 Arrangements Matter
@@
To fill multiple images into the collage prompt, the arrangement of these images could be various. As shown in Figure 1(a), by setting different I, the arrangement of images within the collage prompt is different. Technically, a quadrant-grid collage has 24 potential image arrangements, whereas a nine-grid format can exceed 360,000 possibilities. Our findings in Figure 1(b) indicate that different arrangements can yield varying levels of accuracy, underscoring the importance of optimal image placement within the collage. To determine the most effective configuration for a collage prompt, we utilize a collage predictor. This predictor is designed to predict the expected accuracy of a specific collage arrangement. To facilitate this, it is necessary to first collect an evaluation dataset that contains evaluation outcomes of various collage prompt.
3.2.1 Evaluation Dataset.

### v2

@@ Abstract
Recent advancements in generative AI have suggested that by taking visual prompts, GPT-4V can demonstrate significant proficiency in visual recognition tasks. Despite its impressive capabilities, the financial cost associated with GPT-4V’s inference presents a substantial barrier to its wide use. To address this challenge, we propose a budget-friendly collage prompting task that collages multiple images into a single visual prompt and makes GPT-4V perform visual recognition on several images simultaneously, thereby reducing the cost. We collect a dataset of various collage prompts to assess its performance in GPT-4V’s visual recognition. Our evaluations reveal several key findings: 1) Recognition accuracy varies with different positions in the collage. 2) Grouping images of the same category together leads to better visual recognition results. 3) Incorrect labels often come from adjacent images. These findings highlight the importance of image arrangement within collage prompt. To this end, we construct a benchmark called CollagePrompt, which offers a platform for designing collage prompt to achieve more cost-effective visual recognition with GPT-4V. A baseline method derived from genetic algorithms to optimize collage layouts is proposed and two metrics are introduced to measure the efficiency of the optimized collage prompt. Our benchmark enables researchers to better optimize collage prompts, thus making GPT-4V more cost-effective in visual recognition. The code and data are available at this project page https://collageprompting.github.io/.
CollagePrompt: A Benchmark for Budget-Friendly Visual Recognition with GPT-4V
Siyu Xu1, Yunke Wang1, Daochang Liu2, Bo Du3, Chang Xu1111Corresponding authors., 1School of Computer Science, The University of Sydney, Australia 2School of Physics, Mathematics&Computing, The University of Western Australia, Australia 3School of Computer Science, Institute of Artificial Intelligence, Wuhan University, China. {s.xu,c.xu,yunke.wang}@sydney.edu.au, daochang.liu@uwa.edu.au, dubo@whu.edu.cn
@@ 1 Introduction
With the rapid development of generative AI, various large language models (LLMs) Chang et al. (2023); Zhao et al. (2023); Thirunavukarasu et al. (2023) have emerged as generative tools. Beyond text, these models have expanded their capabilities to include text-to-image generation such as Stable Diffusion Rombach et al. (2022), and text-to-video generation, as seen in Sora Brooks et al. (2024). ChatGPT Brown et al. (2020), as the most well-known LLMs, has shown it can have natural and coherent conversations, making it a powerful tool in daily life and different industry fields. As the latest version of ChatGPT, GPT-4V is a multi-modal LLM capable of processing both text and images. This capability allows it to be applied to a wider range of applications and tasks. There are many technical reports and user studies Li et al. (2023); Lin et al. (2023); Shi et al. (2023); Wen et al. (2023); Yang et al. (2023a); Zhou et al. (2023) about GPT-4V, which conducted thorough evaluations of its capabilities from various aspects.
@@
In Wu et al. (2023c), the visual capabilities of GPT-4V are investigated within the framework of zero-shot visual recognition tasks, such as image and video recognition. The evaluation of visual capabilities is quite straightforward: images and candidate categories are directly fed into GPT-4V for relevance ranking, yielding Top-1 and Top-5 prediction results. Video and point cloud data are uniformly sampled to generate a set of images, which are then processed by GPT-4V for visual recognition. GPT-4V has achieved remarkable performance across various visual recognition tasks, surpassing previous customized solutionsWang et al. (2018b); Dosovitskiy et al. (2020). However, its financial cost associated with its inference can be significant. Specifically, performing image recognition on the ImageNet-1K dataset  Russakovsky et al. (2015) requires approximately \$1 for every 20 images, leading to a total evaluation cost of over \$2,500 for the entire dataset. If we consider the rate limits of maximum API requests per minute, the costs will be even higher. Thus, employing GPT-4V for visual recognition tasks is expensive, and it is meaningful to adopt a more budget-friendly way for GPT-4V’s visual recognition.
@@
Figure 1: Visual recognition of GPT-4V with different prompting ways. (a): Standard Prompt takes one image as visual prompt for each GPT-4V’s run. (b): Collage Prompt concatenates multiple images into one visual prompt and predicts class for all images in each inference. (c): The arrangement of images within collage prompting leads to significantly different results. Green indicates an accurate prediction while red indicates an wrong prediction.
@@
In GPT-4V’s visual recognition, only an image is used as the visual prompt. This standard prompting way fails to fully release the potential capacity of GPT-4V, which is able to process multiple inquiries within one prompt simultaneously. Motivated by this idea, we propose Collage Prompting, a new task of prompting for GPT-4’s visual recognition for budget-friendly inference. In collage prompt, multiple images are collaged into one visual prompt with equal size. GPT-4V is then requested to recognize the class for all images within this prompt. The overall process is shown in Figure 1. Since collage prompting allows for the recognition of multiple images in one run of GPT-4V, it significantly reduces the average cost of visual recognition.
Benchmark. Based on the observation that different arrangements in the collage prompt could lead to rather large variance of accuracy in GPT-4V’s recognition, we further construct a benchmark for arrangement optimization. The benchmark collects a comprehensive dataset that contains various collage prompts, which is then used to either assess the performance of collage prompts or provide a platform to develop algorithms that can optimize collage prompts for more cost-effective recognition. Based on the idea of genetic algorithm, we develop a baseline Learn to Collage (LCP) to optimize the arrangement. In our baseline, the collage prompt is represented as a graph, and a collage predictor is used to estimate the expected accuracy of this collage prompt. LCP is then used to search for the best arrangement in several iterations. Two new metrics are proposed to measure the cost-effectiveness of the developed algorithm.
Contributions. We make three contributions in this paper. First, we propose a budget-friendly prompting approach for GPT-4V. By involving multiple images into a single visual prompt, GPT-4V can process multiple images in one inference run, thus reducing the overall expense greatly. Second, we collect the benchmark dataset of collage prompt. The datasets contain various collage prompts from the ImageNet-1K training set and their accuracy in GPT-4V’s image recognition. This dataset is meaningful for studying the effectiveness of collage prompting. Third, we propose a genetic algorithm-based optimization method for collage arrangement. This approach aims to optimize the arrangement for collage prompts and improve image recognition accuracy within GPT-4V.
@@ 2 Related Works
Exploration of GPT-4V. The state-of-the-art large multi-modal model GPT-4V was firstly launched at September 2023 and has demonstrated its strong visual capability in different fields. Early works Wu et al. (2023d); Yang et al. (2023a) conducted a user study of GPT-4V, where operations were completed by entering prompts on a web interface for GPT-4V. Furthermore, the release of GPT-4V API in November 2023 opened up new opportunities for both academic community and industry to thoroughly evaluate GPT-4V’s performance across various visual benchmarks and provide quantitative data beyond what user studies can offer. GPT-4V’s capacities in multimodal medical diagnosis has been explored in Wu et al. (2023a); Yang et al. (2023b); Deng et al. (2024), where GPT-4V can process different imaging modalities like CT and MRI in medical scene. GPT-4V can also be utilized to operate robots from providing instructions by taking multimodal input in autonomous driving Cui et al. (2024); Han et al. (2024); Wen et al. (2023) and task planning Wake et al. (2023); Wang et al. (2024); Hu et al. (2023). Additionally, GPT-4V has been widely used in advancing video understanding Lin et al. (2023), conducting OCR recognition Shi et al. (2023), acting as an intelligent web agent Zheng et al. (2024), and dealing with each observation data Zhang and Wang (2024).  Wu et al. (2023c) is the first work that considers to adopt extensive quantitative analysis utilizing the established visual benchmarks. However, the evaluation of GPT-4V on visual benchmarks could lead to large expense and it is important to adopt a budget-friendly inference scheme for the evaluation of GPT-4V.
@@
Prompt Engineering in LLMs. Prompt engineering has emerged as a crucial technique for unlocking the potential of pre-trained large language models (LLMs) and vision-language models (VLMs). The concept of prompt engineering was initially explored and popularized in the LLMs Liu et al. (2023); Tonmoy et al. (2024); Chen et al. (2023) and VLMs Wu et al. (2023b); Bahng et al. (2022). The most common prompting way is zero-shot prompting Radford et al. (2019); Cheng et al. (2023), which offers a paradigm shift in leveraging large LLMs. This method significantly reduces the dependency on vast amounts of training data by employing strategically formulated prompts to steer the model towards executing new tasks. While the primary focus in the field has been on creating prompts that can release the potential capacities of LLMs, we focus on developing a prompting approach that prioritizes cost-efficiency.
@@
Figure 2: (a): The workflow of forming the collage prompt from a set of images and related indexes. For a set of images X with two different position indexes I, we can obtain two collage prompts $\textbf{M}_{1}$ and $\textbf{M}_{2}$. Regarding X as the node of a graph, the adjacency matrix of $\textbf{M}_{1}$ and $\textbf{M}_{2}$ can be represented as $\textbf{A}_{1}$ and $\textbf{A}_{2}$. (b): The average accuracy of collage prompts within evaluation datasets using the ‘Worst,’ ‘Average’, and ‘Best’ arrangement.
3 Collage Prompting
GPT-4V has enabled us to perform comprehensive visual recognition. However, each inference performed by GPT-4V incurs a financial cost, which is determined by the number and type of input and generated tokens***https://openai.com/pricing, based on pricing as of March 1, 2024.. Specifically, for image recognition tasks involving images of $512\times 512$ resolution, approximately 5000 tokens are consumed per image. Standard prompting of GPT-4V involves presenting a single image as a visual prompt and processing each image in the dataset individually. With this standard prompting, the expense of evaluating a dataset with 10,000 images could exceed \$500. This method is costly, and a more budget-friendly approach is to process multiple images simultaneously in a single inference run.
Motivated by this idea, we propose Collage Prompting, an efficient alternative to standard visual prompting. Collage Prompting involves concatenating multiple images into a single visual prompt, allowing for simultaneous processing in a single inference run. For example, employing a nine-grid collage prompt can decrease expenses to just $1/9$ of what standard individual image prompting incurs. Moreover, collage prompt not only significantly reduces costs but also processes multiple images with a single API request, thereby reducing server load and inference time. This approach is particularly beneficial for large-scale, high-frequency applications of multimodal foundational models, such as using GPT-4V for image captioning or employing GPT-4o for reading video streams.
@@ 3 Collage Prompting
Preliminary of collage prompt. By assembling $K$ images into one visual prompt, the collage prompt M is designed to be a $\sqrt{K}\times\sqrt{K}$ grid and each grid contains one image. For example, a collage of four images might be presented in a quadrant grid, while nine images could be arranged in a nine-grid format. Supposing we have a set of $K$ images $\textbf{X}=[x_{1},x_{2},...,x_{K}]$ and its related position indexes $\textbf{I}=[i_{1},i_{2},...,i_{K}]$, where $i_{j}$ indicates the position number of image $x_{j}$ in the collage prompt, starting from 00 in the top left corner to $K-1$ in the bottom right corner. Hence, the row position $r$ and column position $c$ of image $x_{j}$ can be specified as $r=\lfloor(i_{j}-1)/\sqrt{K}\rfloor$ and $c=(i_{j}-1)\mod\sqrt{K}$ respectively. While collage prompt M has the same size as the standard prompt, GPT-4V can thus take this collage prompt as input and generate the predicted class for all images within the collage prompt. Regarding the collage prompt M as graph $\textbf{M}=(\textbf{A},\textbf{F})$ with $K$ nodes, each node $f_{i}\in\mathbb{R}^{l}$ denotes the feature of $x_{i}$. The adjacency matrix $\textbf{A}\in\mathbb{R}^{K\times K}$ suggests the relative positions of X, where two adjacent images are considered to have an undirected edge. For two images $x_{p}$ and $x_{q}$ that satisfies either $|c_{p}-c_{q}|=1$ when $r_{p}=r_{q}$ or $|r_{p}-r_{q}|=1$ when $c_{p}=c_{q}$, we consider these two images have an edge and set $\textbf{A}[p,q]=\textbf{A}[q,p]=1$. The workflow of representing the collage prompt as a graph is illustrated in Figure 2(a).
@@
Arrangements matter. To fill multiple images into the collage prompt, the arrangement of these images could be various. As shown in Figure 2(a), by setting different I, the arrangement of images within the collage prompt is different. Technically, a quadrant-grid collage has $24(4!)$ potential image arrangements, whereas a nine-grid format can exceed $360,000(9!)$ possibilities. Our findings in Figure 2(b) indicate that different arrangements can yield varying levels of accuracy, underscoring the importance of image arrangement. Therefore, we hope to find better arrangement within the collage to minimize the accuracy loss of GPT-4V.
4 A Benchmark of Collage Prompting
The arrangement of images within a collage prompt significantly impacts the overall recognition accuracy of GPT-4V. For any given set of images forming a collage, there should exist one or more optimal arrangements that maximize overall recognition accuracy. Thus, we conduct a benchmark study of collage prompting with two primary objectives: 1) to study the effect of different arrangements on GPT-4V’s recognition accuracy and 2) to provide a benchmark for developing algorithms to optimize collage arrangements. In this section, we first present a comprehensive collage prompt dataset to assess the performance of various collage prompts. Three key observations suggest that there is a need to conduct arrangement optimization. Then, we propose a baseline method to learn the layout of the collage and two metrics that reflect the cost-effective trait of GPT-4V’s visual recognition are proposed.
(a) "0: ‘great white shark’, 1: ‘stingray’, 2: ‘great white shark’, 3: ‘great white shark’"
(b) "0: ‘great white shark’, 1: ‘stingray’, 2: ‘great white shark’, 3: ‘stingray’"
(c) "0: ‘tractor’, 1: ‘automated teller machine’, 2: ‘red-breasted merganser’, 3: ‘dive’"
(d) "0: ‘automated teller machine’, 1: ‘tractor’, 2: ‘dock’, 3: ‘red-breasted merganser’"
Figure 3: (a) and (b) demonstrate the effect of category clustering, where placing images of the same category together increases overall recognition accuracy. (c) and (d) illustrate localization errors, where GPT-4V predicts the correct labels but outputs them to incorrect positions in the collage.
Figure 4: Average Prediction Accuracy by Position.
4.1 Dataset